OpenStudy (anonymous):

what is 5(3/10)²-3(3/10)+2

OpenStudy (jamesj):

First thing I'd do is factor out 1/10^2. Then your expression is equal to \[ \frac{1}{10^2} [ 5(3^2) - 3(3)(10) + 2(10^2) ] \] So far so good?

OpenStudy (anonymous):

wao thats complicated

OpenStudy (jamesj):

Well, now the expression in square brackets is \[ 45 - 90 + 200 = 155 \] Yes?

OpenStudy (anonymous):

ok yes but... can u show me why its in brackets, and isnt that like reverse simplifying?

OpenStudy (anonymous):

oh wait nvm i get it

OpenStudy (jamesj):

I just put the numbers in ( ) type brackets to emphasize we're multiplying them. The square brackets [ ] I used so there wouldn't be so many round brackets.

OpenStudy (jamesj):

Now that means the entire expression is \[ \frac{155}{10^2} = \frac{155}{100} \] Now you just need to simplify that.

OpenStudy (anonymous):

yeah ok... not hard at all

OpenStudy (anonymous):

is that... 1 ?

OpenStudy (jamesj):

Is \[ \frac{155}{100} = 1 \ \ \ ??!? \] No way.

OpenStudy (anonymous):

oh wait nvm umm 15.5? no way 1.55?

OpenStudy (anonymous):

i think its 1.55?

OpenStudy (jamesj):

your final answer is ... yes, correct.

OpenStudy (anonymous):

wooo thanks